OpenStudy (anonymous):

write 3sin2xsin3x as a sum or difference

OpenStudy (anonymous):

sum formula for sin/cos: sin2x=2sinxcosx, and cos2x=cos^2x-sin^2x notice that sin3x=sin(x+2x)=sinxcos2x+sin2xcosx=sinx(cos^2x-sin^2x)+2sinxcosx*cosx NOTE: sin^2x is my notation for (sinx)^2 and similarly for cosine. so in your original problem, 3(sin2x)(sin3x) =3 (2sinx cosx) (sinx (cos^2x - sin^2x) + 2sinx cosx cosx) =6 sinx cosx (sinx cos^2x-sin^3x + 2 sinx cos^2x) =6 sin^2x cos^3x - 6sin^4x cosx + 12 sin^2x cos^3x